Hydrogen gas has been suggested as a clean fuel because it produces only water vapor when it burns. If the reaction has 77.6% yield, what mass of hydrogen forms 682 kg of water, enter answer in scientific notation
1
Expert's answer
2019-09-30T05:14:40-0400

2H2+O2=2H2O

682 kg = 682000 g.

(682000/36 )x 4= 75777.7776 g.

75,777.7776/ 0.776=97651.7753g

97,651.7753-75,777.7776= 21873.9977g= 21.8739977 kg.